OpenStudy (anonymous):

ANOTHER DIFFICULT ONE!...for me.. Let f(x)=X^2-100 . The average rate of change from x = -10 to x = 10 equals _____.

OpenStudy (anonymous):

A. 100 B. 200 C. 0 D. -100

OpenStudy (paxpolaris):

\[\LARGE ={f(10)-f(-10) \over 10-(-10)}\]

OpenStudy (anonymous):

0?

OpenStudy (paxpolaris):

yes

OpenStudy (anonymous):

THANKS PAX!

OpenStudy (anonymous):

imma fan :)

OpenStudy (anonymous):

wouldn't that be undefined?

OpenStudy (anonymous):

or we take a limit?

OpenStudy (paxpolaris):

\[{0 \over 20} = 0\]

OpenStudy (anonymous):

aaa ok

OpenStudy (anonymous):

I was thinking something else

OpenStudy (anonymous):

Let f(x)=1/x be the reciprocal function. The average rate of change from x = 0.01 to x = 0.1 equals _____.

OpenStudy (anonymous):

do u guys think that this is -10

OpenStudy (anonymous):

@PaxPolaris @timo86m ?

OpenStudy (anonymous):

i believe its -10

OpenStudy (anonymous):

its 0 I am with pax

OpenStudy (anonymous):

no for the new question i just asked.

OpenStudy (anonymous):

oh just follow pax's example (f(xf)-f(xi))/ ( xf -xi )

OpenStudy (anonymous):

so then its still 0 ?

OpenStudy (anonymous):

o.O

OpenStudy (paxpolaris):

\[\LARGE ={f(0.1)-f(.01) \over .1-(.01)}={-90 \over 0.09}=?\] not -10

OpenStudy (anonymous):

-1000!

OpenStudy (anonymous):

I GUESS ITS ALL ABOUT HAVING THE RIGHT FUNCTION FOR EACH QUESTION, BUT ITS FIGURING OUT THAT FUNCTION THAT IS HARD.

OpenStudy (anonymous):

-1000

OpenStudy (anonymous):

what does it mean when it says the rate of change?

OpenStudy (paxpolaris):

average rate of change is the SLOPE of the line passing through the two points \[\LARGE ={y_2-y_1 \over x_2-x_1}\]

OpenStudy (anonymous):

so dont you have to have cordinants to find that?

OpenStudy (paxpolaris):

you have the x co-ords (eg -10 & 10 in the first question) & you can calculate y=f(x)

OpenStudy (anonymous):

because for another hard question i had it says f(x)=1/2x+5 then it states The average rate of change from x = 17 to x = 43 equals but didnt you say y2-y 1/x2-x1

OpenStudy (anonymous):

all i have is the x cordinants

OpenStudy (paxpolaris):

you have the equation to find y .... or f(x) if you prefer

OpenStudy (anonymous):

ooo so y=1/2x+5

OpenStudy (anonymous):

how do i get the y cordinants to look like the x cordinants ?

OpenStudy (paxpolaris):

\[\LARGE ={f(43)-f(17) \over 43-17}\] \[f(43) = {1 \over 2(43)+5}\]

OpenStudy (anonymous):

1/91?

OpenStudy (paxpolaris):

yes.. keep going ... find the final answer

OpenStudy (anonymous):

ahh can u give me a hint on what to do next :(

OpenStudy (paxpolaris):

whats f(17)

OpenStudy (anonymous):

1/39

OpenStudy (anonymous):

?

OpenStudy (paxpolaris):

yes

OpenStudy (paxpolaris):

i did read the equation right ? right? you meant : \[\large f(x) = {1 \over 2x+ 5}\]

OpenStudy (anonymous):

uhm f(x)=1/2 x <- the x is on the side of the fraction then +5

OpenStudy (anonymous):